# GANQ: GPU-Adaptive Non-Uniform Quantization for Large Language Models

### v2

@@ 1 Introduction
Quantization is a promising solution to reduce inference costs for LLMs. For example, 4-bit weight quantization can reduce memory usage for model loading by nearly 75% compared to FP16. In general, quantization techniques are categorized into quantization-aware training (QAT) and post-training quantization (PTQ). QAT integrates quantization into the training process to achieve higher accuracy but is computationally expensive, often requiring extensive samples and significant GPU hours (Liu et al., 2023). This makes QAT impractical for large models. In contrast, PTQ is a cost-effective alternative that applies quantization after training, making it the preferred choice for LLMs (Nagel et al., 2020; Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023; Kim et al., 2023; Lin et al., 2024; Shao et al., 2024; Ma et al., 2024; Li et al., 2024b). Among PTQ methods, weight-only quantization, which uses low-precision weights while retaining high-precision activations, has become a particularly attractive approach. By reducing memory traffic and alleviating memory-bound bottlenecks, weight-only quantization accelerates inference (Kim et al., 2023). Additionally, compared to weight-activation quantization, it avoids significant accuracy degradation by preserving the precision of activations, ensuring better model performance.
@@
To address these issues, we propose GANQ (GPU-Adaptive Non-Uniform Quantization), a layer-wise post-training non-uniform quantization framework optimized for lookup table (LUT)-based mpGEMM. In LUT-based mpGEMM (see the right part of Figure 1(a)), complex computations are replaced with simple table lookups, supported by several GPU kernels (Kim et al., 2023; Mo et al., 2024). The primary challenge then becomes how to determine effective low-bit representations for the LUTs. Existing non-uniform quantization methods often rely on heuristic-based approaches, such as manually designed mappings (e.g., power-exponent functions (Yvinec et al., 2023)) or clustering-based methods with heuristic distance metrics (Han et al., 2015; Xu et al., 2018; Kim et al., 2023). While these methods may achieve good results in specific cases, their heuristic nature limits generalization and theoretical grounding. In contrast, GANQ introduces a principled optimization model for layer-wise LUT-based non-uniform quantization, formulated as a mixed-integer quadratic programming problem. This model minimizes the the discrepancy between the outputs of the quantized and original layers, thereby preserving accuracy. To efficiently address this complex model, GANQ utilizes its decomposable structure to divide the original optimization task into multiple independent one-dimensional subproblems, which can be processed in parallel using GPU acceleration to achieve substantial computational efficiency. Besides, GANQ employs an alternating direction optimization framework that capitalizes on the splittable structure of decision variables, effectively reducing quantization error.
In addition, although GANQ is designed as a base quantization method, it is fully compatible with current techniques for handling outliers, such as splitting weights into sparse components (to address outliers) and quantized components (Dettmers et al., 2023; Kim et al., 2023), thereby enabling further performance enhancements.
We evaluate GANQ extensively across various model families and sizes on language modeling tasks. The results show that GANQ consistently outperforms previous methods in quantization performance. Moreover, GANQ is highly resource-efficient and easy to implement. For instance, GANQ processes the LLaMA-2-7B model on a single NVIDIA RTX 4090 GPU in approximately one hour using only 128 samples. Additionally, our deployed models on the NVIDIA RTX 4090 GPU achieve significant latency improvements, with speedups of up to 2.57$\times$ compared to the FP16 baseline. These results highlight the effectiveness of GANQ in both quantization quality and inference efficiency.
@@ 2 Related Work
Quantization for LLMs. Quantization reduces the bit-precision of neural networks, resulting in smaller models and faster inference. It has become a key direction for compressing LLMs given their growing size and inference costs. Current quantization methods for LLMs are broadly categorized into QAT (Liu et al., 2023) and PTQ (Nagel et al., 2020; Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023; Kim et al., 2023; Lin et al., 2024; Shao et al., 2024; Ma et al., 2024; Li et al., 2024b). QAT integrates quantization into the training process, preserving high performance but incurring prohibitive training costs, making it impractical for LLMs. In contrast, PTQ applies quantization to pretrained models, requiring only a small subset of data and modest computational resources, making it particularly appealing for LLMs. PTQ methods can be further classified into wight-only quantization and weight-activation quantization.
@@
Weight-only quantization focuses on compressing model weights into low-bit formats. For example, GPTQ (Frantar et al., 2022) utilizes the optimal brain surgeon framework (Hassibi & Stork, 1992) for quantization and reconstruction. OminiQuant (Shao et al., 2024) introduces learnable parameters to determine quantization factors (e.g., scale and zero-point), while AffineQuant (Ma et al., 2024) extends this idea by incorporating a learnable matrix to preprocess weights before quantization.
Weight-activation quantization compresses both weights and activations, often addressing their quantization jointly. For example, SmoothQuant (Xiao et al., 2023) shifts quantization difficulty from activations to weights using manually designed scaling factors. Similarly, SVDQuant (Li et al., 2024b) applies this approach while further decomposing weights into low-rank and quantized components.
@@
While weight-activation quantization can offer broader compression, studies (Kim et al., 2023; Lin et al., 2024) have shown that LLM inference, especially during generation, is heavily memory-bound, with weight access dominating activation access by orders of magnitude. Consequently, weight-only quantization is more effective for on-device deployment of LLMs. In this work, we focus on weight-only PTQ for its efficiency and suitability for LLMs.
@@
Outlier Mitigation. Due to the widely used uniform quantization mapping and the inherent non-uniform distribution of LLM weights, a key challenge is the presence of outliers. These outliers unnecessarily expand the quantization range (see Figure 1(b)), comprising quantization performance. Recent methods have been proposed to address this issue. For example, SpQR (Dettmers et al., 2023) and SqueezeLLM (Kim et al., 2023) retain outliers in sparse matrices while applying quantization to the remaining weights to mitigate their impact on overall performance. AWQ (Lin et al., 2024) independently quantizes the channel-wise salient weights to improve performance, and SVDQuant (Li et al., 2024b), as mentioned, decomposes weights into low-rank and quantized components. While these methods effectively handle outliers and enhance quantization performance, they often introduce additional computational overhead during inference. For instance, SpQR and SqueezeLLM require both mpGEMM and sparse matrix multiplication, whereas SVDQuant adds an extra low-rank computation branch.
@@
Non-Uniform Quantization. The non-uniform distribution of weights in LLMs highlights the importance of non-uniform quantization. However, existing non-uniform quantization methods often rely on heuristic-based approaches, limiting their generalization and theoretical grounding. NUPES (Yvinec et al., 2023) replaces uniform quantization with power-exponent functions and employs gradient-based optimization to learn the exponent parameter. Other methods focus on identifying shared weights, thereby forming a codebook, which is suitable for LUT-based mpGEMM. For example, Han et al. (2015) apply $k$-means clustering to minimize the Euclidean distance between weights and centroids in convolutional neural networks (CNNs), while Xu et al. (2018) extend this approach by using a loss-based metric for $k$-means clustering in CNNs. For LLMs, SqueezeLLM (Kim et al., 2023) adapts this idea by leveraging sensitivity-based $k$-means clustering, where the sensitivity metric measures measures the extent to which the model is perturbed after quantization. To mitigate the computational expense of this calculation, SqueezeLLM approximates the required Hessian matrix using the diagonal elements of the Fisher information matrix (Fisher, 1925).
@@ 3.1 Optimization Model for Non-uniform Quantization
Consider a linear layer with weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ and input activation $\mathbf{X}\in\mathbb{R}^{n\times p}$. As shown in the right part of Figure 1(a), LUT-based quantization aims to compress $\mathbf{W}$ by representing its elements using a codebook. Specifically, the elements of the $i$-th channel in the quantized weight matrix $\mathbf{\widetilde{W}}$ are selected from the codebook $\mathbf{T}_{i}=\{t_{i,0},t_{i,1},\dots,t_{i,2^{N}-1}\}$, where $N$ is the bit-width of the quantization (e.g., 3 or 4 bits). Thus, each element $\mathbf{\widetilde{W}}_{i,j}$ satisfies $\mathbf{\widetilde{W}}_{i,j}\in\mathbf{T}_{i}$.
@@
$m=n=2048$ (e.g., $\mathbf{W}_{\{k,q,v,o\}}$ in OPT-1.3B) $100.00\%$ $25.10\%$ $25.78\%$
$m=n=4096$ (e.g., $\mathbf{W}_{\{k,q,v,o\}}$ in LLaMA-2-7B) $100.00\%$ $25.05\%$ $25.39\%$
$m=n=8192$ (e.g., $\mathbf{W}_{\{k,q,v,o\}}$ in LLaMA-2-70B) $100.00\%$ $25.02\%$ $25.20\%$
@@
The mixed-integer structure of $\mathbf{S}_{i}$ introduces significant combinatorial complexity, and the bilinear interaction between $\mathbf{S}_{i}$ and $\mathbf{T}_{i}$ in the objective further compounds the computational challenge, rendering the problem inherently non-convex and non-smooth. These factors pose serious difficulties for off-the-shelf solvers, especially in large-scale settings with high-dimensional weight matrices and input activations. In response, we develop a specialized, GPU-adaptive approach tailored to navigate this complex search space while scaling to practical problem sizes.
@@ 3.2 GPU-Adaptive Non-uniform Quantization Method
The $\mathbf{T}_{i}$-subproblem in (4) is an unconstrained quadratic program and admits a closed form solution given by:
@@
If $\mathbf{X}\mathbf{X}^{\top}$ is not positive definite, which is rare but can occur in cases like the fc2 layer of OPT models, we can add $\lambda\mathbf{I}$ ($\lambda>0$) to guarantee positive definiteness before Cholesky decomposition. Specifically, for any non-zero vector $\mathbf{v}$, $\mathbf{v}^{\top}(\mathbf{X}\mathbf{X}^{\top}+\lambda\mathbf{I})\mathbf{v}=\|\mathbf{X}^{\top}\mathbf{v}\|^{2}+\lambda\|\mathbf{v}\|^{2}>0$.
@@
      $\mathbf{r}=(\mathbf{W}_{:,j:}-\mathbf{T}^{k}\mathbf{S}_{:,:,j:}^{k+1})\mathbf{L}_{j:,j}$ update residual
@@ 3.3 Compatibility with Outlier-Handling Techniques
GANQ provides a foundational framework for LUT-based non-uniform quantization and is inherently compatible with existing techniques for handling outliers in weight matrices. Among these techniques, a widely adopted approach involves splitting the weight matrix into a sparse matrix for outliers and a quantized matrix for the remaining weights. For example, SpQR (Dettmers et al., 2023) and SqueezeLLM (Kim et al., 2023) extract outliers into a separate sparse matrix to mitigate their impact on the quantization process.
In our framework, the weight matrix $\mathbf{W}$ can similarly be decomposed into a sparse component $\mathbf{W}_{\text{sparse}}$, containing extracted outliers, and a dense component $\mathbf{W}_{\text{dense}}$, processed through GANQ. The method for extracting outliers is detailed in Appendix A. This decomposition reduces quantization range, thereby enhancing the quantization performance.
@@ 4.1 Settings
Models. We comprehensively evaluate GANQ on a range of models, including OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), and LLaMA-3 (Meta AI, 2024) families. Specifically, we assess its performance across OPT-125M, OPT-350M, OPT-1.3B, OPT-2.7B, OPT-6.7B, LLaMA-7B, LLaMA-2-7B, and LLaMA-3-8B models.
Evaluation. Following prior work (Frantar et al., 2022; Shao et al., 2024; Ma et al., 2024; Kim et al., 2023), we evaluate the quantized models by reporting perplexity on language generation tasks, specifically using the WikiText-2 (Merity et al., 2016), C4 (Raffel et al., 2020), and PTB (Marcus et al., 1994) datasets. Additionally, we assess accuracy on zero-shot tasks, including ARC Easy, ARC Challenge (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), RTE (Wang et al., 2018), and HellaSwag (Zellers et al., 2019), facilitated by the LM Harness library (Gao et al., 2021).
Baselines. For basic weight-only quantization, we compare GANQ with standard round-to-nearest uniform quantization (RTN), GPTQ (Frantar et al., 2022), and OminiQuant (Shao et al., 2024). For weight-only quantization with outlier handling, we compare with GPTQ, OminiQuant, and AWQ (Lin et al., 2024), each using a group size of 128, as well as SqueezeLLM (Kim et al., 2023).
Setup. We implement GANQ using PyTorch (Paszke et al., 2019) and utilize the HuggingFace Transformers library (Wolf, 2019) for model and dataset management. All experiments are conducted on a single NVIDIA RTX 4090 GPU. For calibration data, we follow the methodology outlined in previous works (Frantar et al., 2022; Shao et al., 2024; Kim et al., 2023). Specifically, we use 32 sequences for OPT models and 128 sequences for LLaMA models. Each sequence consists of 2048 tokens, sampled from the first shard of the C4 dataset.
Latency Profiling. Following prior works and associated quantized model inference kernels (GPTQ-for-LLaMa, 2023; Kim et al., 2023), we evaluate latency and peak memory usage while generating 1024 tokens on a single NVIDIA RTX 4090 GPU using the Torch CUDA profiler.

### v3

@@ 1 Introduction
Quantization is a promising solution to reduce inference costs for LLMs. For example, 4-bit weight quantization can reduce memory usage for model loading by nearly 75% compared to FP16. In general, quantization techniques are categorized into quantization-aware training (QAT) and post-training quantization (PTQ). QAT integrates quantization into the training process to achieve higher accuracy but is computationally expensive, often requiring extensive samples and significant GPU hours (Liu et al., 2024). This makes QAT impractical for large models. In contrast, PTQ is a cost-effective alternative that applies quantization after training, making it the preferred choice for LLMs (Nagel et al., 2020; Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023; Kim et al., 2024; Lin et al., 2024; Shao et al., 2024; Ma et al., 2024; Li et al., 2024b). Among PTQ methods, weight-only quantization, which uses low-precision weights while retaining high-precision activations, has become a particularly attractive approach. By reducing memory traffic and alleviating memory-bound bottlenecks, weight-only quantization accelerates inference (Kim et al., 2024; Lin et al., 2024). Additionally, compared to weight-activation quantization, it avoids significant accuracy degradation by preserving the precision of activations, ensuring better model performance.
@@
To address these issues, we propose GANQ (GPU-Adaptive Non-Uniform Quantization), a layer-wise post-training non-uniform quantization framework optimized for lookup table (LUT)-based mpGEMM. In LUT-based mpGEMM (see the right part of Figure 1(a)), complex computations are replaced with simple table lookups, supported by several GPU kernels (Kim et al., 2024; Mo et al., 2024; Guo et al., 2024). The primary challenge then becomes how to determine effective low-bit representations for the LUTs. Existing non-uniform quantization methods often rely on heuristic-based approaches, such as manually designed mappings (e.g., power-exponent functions (Yvinec et al., 2023)) or clustering-based methods with heuristic distance metrics (Han et al., 2015; Xu et al., 2018; Kim et al., 2024). While these methods may achieve good results in specific cases, their heuristic nature limits generalization and theoretical grounding. In contrast, GANQ introduces a principled optimization model for layer-wise LUT-based non-uniform quantization, formulated as a mixed-integer quadratic programming problem. This model minimizes the discrepancy between the outputs of the quantized and original layers, thereby preserving accuracy. To efficiently address this complex model, GANQ utilizes its decomposable structure to divide the original optimization task into multiple independent one-dimensional subproblems, which can be processed in parallel using GPU acceleration to achieve substantial computational efficiency. Besides, GANQ employs an alternating direction optimization framework that capitalizes on the splittable structure of decision variables, effectively reducing quantization error.
In addition, although GANQ is designed as a base quantization method, it is fully compatible with current techniques for handling outliers, such as splitting weights into sparse components (to address outliers) and quantized components (Dettmers et al., 2023; Kim et al., 2024), thereby enabling further performance enhancements.
We evaluate GANQ extensively across various model families and sizes on language modeling tasks. The results show that GANQ consistently outperforms previous methods in quantization performance. Moreover, GANQ is highly resource-efficient and easy to implement. For instance, GANQ processes the LLaMA-2-7B model on a single NVIDIA RTX 4090 GPU in approximately one hour, using only 128 samples, each containing 2,048 tokens. Furthermore, our deployed models on the NVIDIA RTX 4090 GPU achieve up to 2.57$\times$ speedups over the FP16 baseline by leveraging LUT-based inference kernels (Kim et al., 2024). These results highlight the effectiveness of GANQ in both quantization quality and inference efficiency.
@@ 2 Related Work
Quantization for LLMs. Quantization reduces the bit-precision of neural networks, resulting in smaller models and faster inference. It has become a key direction for compressing LLMs given their growing size and inference costs. Current quantization methods for LLMs are broadly categorized into QAT (Liu et al., 2024) and PTQ (Nagel et al., 2020; Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023; Kim et al., 2024; Lin et al., 2024; Shao et al., 2024; Ma et al., 2024; Li et al., 2024b). QAT integrates quantization into the training process, preserving high performance but incurring prohibitive training costs, making it impractical for LLMs. In contrast, PTQ applies quantization to pretrained models, requiring only a small subset of data and modest computational resources, making it particularly appealing for LLMs. PTQ methods can be further classified into wight-only quantization and weight-activation quantization.
@@
Weight-only quantization focuses on compressing model weights into low-bit formats. For example, GPTQ (Frantar et al., 2022) utilizes the optimal brain surgeon framework (Hassibi & Stork, 1992) for quantization and reconstruction. OmniQuant (Shao et al., 2024) introduces learnable parameters to determine quantization factors (e.g., scale and zero-point), while AffineQuant (Ma et al., 2024) extends this idea by incorporating a learnable matrix to preprocess weights before quantization. Weight-activation quantization compresses both weights and activations, often addressing their quantization jointly. For example, SmoothQuant (Xiao et al., 2023) shifts quantization difficulty from activations to weights using manually designed scaling factors. Similarly, SVDQuant (Li et al., 2024b) applies this approach while further decomposing weights into low-rank and quantized components.
@@
While weight-activation quantization can offer broader compression, studies (Kim et al., 2024; Lin et al., 2024) have shown that LLM inference, especially during generation, is heavily memory-bound, with weight access dominating activation access by orders of magnitude. Consequently, weight-only quantization is more effective for on-device deployment of LLMs. In this work, we focus on weight-only PTQ for its efficiency and suitability for LLMs.
@@
Outlier Mitigation. Due to the widely used uniform quantization mapping and the inherent non-uniform distribution of LLM weights, a key challenge is the presence of outliers. These outliers unnecessarily expand the quantization range (see Figure 1(b)), comprising quantization performance. Recent methods have been proposed to address this issue. For example, SpQR (Dettmers et al., 2023) and SqueezeLLM (Kim et al., 2024) retain outliers in sparse matrices while applying quantization to the remaining weights to mitigate their impact on overall performance. AWQ (Lin et al., 2024) independently quantizes the channel-wise salient weights to improve performance, and SVDQuant (Li et al., 2024b), as mentioned, decomposes weights into low-rank and quantized components. While these methods effectively handle outliers and enhance quantization performance, they often introduce additional computational overhead during inference. For instance, SpQR and SqueezeLLM require both mpGEMM and sparse matrix multiplication, whereas SVDQuant adds an extra low-rank computation branch.
@@
LUT-based Inference Kernel. Low-bit quantized LLMs depend on mpGEMM for efficient inference. This operation, which involves multiplying low-precision weights with higher-precision activations, presents a critical computational challenge. Current hardware lacks native support for mpGEMM, compelling existing implementations to adopt dequantization-based workarounds (Frantar et al., 2022; Lin et al., 2024; Shao et al., 2024; Ma et al., 2024; Li et al., 2024b). LUT-based methods offer a compelling alternative by eliminating dequantization overhead. Through efficient substitution of arithmetic operations with table lookups (Kim et al., 2024; Mo et al., 2024; Guo et al., 2024), these approaches demonstrate particular suitability for mpGEMM acceleration.
Unlike previous work focused on kernel-level optimizations, our research instead targets fundamental improvements in LUT-based quantization accuracy and compatibility with existing kernels.
Non-Uniform Quantization. The non-uniform distribution of weights in LLMs highlights the importance of non-uniform quantization. However, existing non-uniform quantization methods often rely on heuristic-based approaches, limiting their generalization and theoretical grounding. NUPES (Yvinec et al., 2023) replaces uniform quantization with power-exponent functions and employs gradient-based optimization to learn the exponent parameter. Other methods focus on identifying shared weights, thereby forming a codebook, which is suitable for LUT-based mpGEMM. For example, Han et al. (2015) apply $k$-means clustering to minimize the Euclidean distance between weights and centroids in convolutional neural networks (CNNs), while Xu et al. (2018) extend this approach by using a loss-based metric for $k$-means clustering in CNNs. For LLMs, SqueezeLLM (Kim et al., 2024) adapts this idea by leveraging sensitivity-based $k$-means clustering, where the sensitivity metric measures the extent to which the model is perturbed after quantization. To mitigate the computational expense of this calculation, SqueezeLLM approximates the required Hessian matrix using the diagonal elements of the Fisher information matrix (Fisher, 1925).
@@ 3.1 Optimization Model for Non-uniform Quantization
Consider a linear layer with weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ and input activation $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n$ represents the input hidden dimension, $m$ the output dimension, and $p=b\times s$ accounts for batched processing of $b$ sequences each of length $s$. As shown in the right part of Figure 1(a), LUT-based quantization aims to compress $\mathbf{W}$ by representing its elements using a codebook. Specifically, the elements of the $i$-th channel in the quantized weight matrix $\mathbf{\widetilde{W}}$ are selected from the codebook $\mathbf{T}_{i}=\{t_{i,0},t_{i,1},\dots,t_{i,2^{N}-1}\}$, where $N$ is the bit-width of the quantization (e.g., 3 or 4 bits). Thus, each element $\mathbf{\widetilde{W}}_{i,j}$ satisfies $\mathbf{\widetilde{W}}_{i,j}\in\mathbf{T}_{i}$.
@@
$m=n=2048$ (e.g., $\mathbf{W}_{q}$ in OPT-1.3B) $100.00\%$ $25.10\%$ $25.78\%$
$m=n=4096$ (e.g., $\mathbf{W}_{q}$ in LLaMA-2-7B) $100.00\%$ $25.05\%$ $25.39\%$
$m=n=8192$ (e.g., $\mathbf{W}_{q}$ in LLaMA-2-70B) $100.00\%$ $25.02\%$ $25.20\%$
@@
The mixed-integer structure of $\mathbf{S}_{i}$ introduces significant combinatorial complexity, and the bilinear interaction between $\mathbf{S}_{i}$ and $\mathbf{T}_{i}$ in the objective further compounds the computational challenge, rendering the problem inherently non-convex and non-smooth. These factors pose serious difficulties for off-the-shelf solvers (Gurobi Optimization, 2025; IBM, 2025), especially in large-scale settings with high-dimensional weight matrices and input activations. In response, we develop a specialized, GPU-adaptive approach tailored to navigate this complex search space while scaling to practical problem sizes.
@@ 3.2 GPU-Adaptive Non-Uniform Quantization Method
The $\mathbf{T}_{i}$-subproblem in (4) is an unconstrained quadratic program that admits a closed-form solution. Specifically, consider the objective function of $\mathbf{T}_{i}$-subproblem:
$f(\mathbf{T}_{i})=\|\mathbf{W}_{i}\mathbf{X}-\mathbf{T}_{i}\mathbf{S}_{i}^{k+1}\mathbf{X}\|^{2}.$ (5)
To minimize this function, we apply the first-order optimality condition. Taking the gradient of $f(\mathbf{T}_{i})$ with respect to $\mathbf{T}_{i}$ and setting it to zero yields:
$\nabla f(\mathbf{T}_{i})=2(\mathbf{W}_{i}\mathbf{X}-\mathbf{T}_{i}\mathbf{S}_{i}^{k+1}\mathbf{X})\mathbf{X}^{\top}(\mathbf{S}_{i}^{k+1})^{\top}=\mathbf{0}.$ (6)
Solving this matrix equation gives the closed-form update for $\mathbf{T}_{i}$:
@@
If $\mathbf{X}\mathbf{X}^{\top}$ is not positive definite, which is rare but can occur in cases like the fc2 layer of OPT models, we can add $\lambda\mathbf{I}$ ($\lambda>0$) to guarantee positive definiteness before Cholesky decomposition. Specifically, for any non-zero vector $\mathbf{v}$, adjusted matrix satisfies $\mathbf{v}^{\top}(\mathbf{X}\mathbf{X}^{\top}+\lambda\mathbf{I})\mathbf{v}=\|\mathbf{X}^{\top}\mathbf{v}\|^{2}+\lambda\|\mathbf{v}\|^{2}>0$. This preconditioning is a standard technique in numerical linear algebra.
Remark 3.1 describes a basic strategy to ensure positive definiteness by augmenting the matrix $\mathbf{X}\mathbf{X}^{\top}$ with a scaled identity matrix. However, selecting an appropriate $\lambda$ manually can be cumbersome and suboptimal. In practice, we adopt an adaptive preconditioning approach that enforces diagonal dominance, inherently ensuring positive definiteness without requiring manual hyperparameter tuning. Details are provided in Appendix A.
@@
      $\mathbf{r}=(\mathbf{W}_{:,j:}-\mathbf{T}^{k}\mathbf{S}_{:,:,j:}^{k+1})\mathbf{L}_{j:,j-1}$ update residual
@@ 3.3 Compatibility with Outlier-Handling Techniques
GANQ provides a foundational framework for LUT-based non-uniform quantization and is inherently compatible with existing techniques for handling outliers in weight matrices. Among these techniques, a widely adopted approach involves splitting the weight matrix into a sparse matrix for outliers and a quantized matrix for the remaining weights. For example, SpQR (Dettmers et al., 2023) and SqueezeLLM (Kim et al., 2024) extract outliers into a separate sparse matrix to mitigate their impact on the quantization process.
In our framework, the weight matrix $\mathbf{W}$ can similarly be decomposed into a sparse component $\mathbf{W}_{\text{sparse}}$, containing extracted outliers, and a dense component $\mathbf{W}_{\text{dense}}$, processed through GANQ. Appendix B details our outlier extraction method, which identifies outliers using a small ratio threshold $r$ (e.g., $r=0.5\%$ of total parameters) while preserving the remaining weights for quantization. This decomposition reduces quantization range, thereby enhancing the quantization performance.
@@ 4.1 Settings
Models. We comprehensively evaluate GANQ on a range of models, including OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), LLaMA-3 (Meta AI, 2024a), and LLaMA-3.2 (Meta AI, 2024b) model families. Specifically, we assess its performance across OPT-125M, OPT-350M, OPT-1.3B, OPT-2.7B, OPT-6.7B, LLaMA-7B, LLaMA-2-7B, LLaMA-3-8B, LLaMA-3.2-1B, LLaMA-3.2-3B, LLaMA-3.2-1B-Instruct, and LLaMA-3.2-3B-Instruct models.
Evaluation. Following prior work (Frantar et al., 2022; Shao et al., 2024; Ma et al., 2024; Kim et al., 2024), we evaluate the quantized models by reporting perplexity on language datasets, specifically using the WikiText-2 (Merity et al., 2017), C4 (Raffel et al., 2020), and PTB (Marcus et al., 1994) datasets. Consistent with established practice, we use a sequence length of 2,048 across all models. Additionally, we assess accuracy on zero-shot tasks, including ARC Easy, ARC Challenge (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), RTE (Wang et al., 2018), HellaSwag (Zellers et al., 2019), and GSM8K (Cobbe et al., 2021), facilitated by the LM Harness library (Gao et al., 2021). Long-context capabilities are evaluated using LongBench (Bai et al., 2024) under its standard protocol.
Baselines. For basic weight-only quantization, we compare GANQ with standard round-to-nearest uniform quantization (RTN), GPTQ (Frantar et al., 2022), and OmniQuant (Shao et al., 2024). For weight-only quantization with outlier handling, we compare with GPTQ, OmniQuant, and AWQ (Lin et al., 2024), each using a group size of 128, as well as SqueezeLLM (Kim et al., 2024).
Setup. We implement GANQ using the PyTorch (Paszke et al., 2019) and utilize the HuggingFace Transformers library (Wolf, 2019) for model and dataset management. Our implementation is publicly available [^1] The code is available at https://github.com/Evans-Z/GANQ. All experiments are conducted on a single NVIDIA RTX 4090 GPU. For calibration data, we follow the methodology outlined in previous works (Frantar et al., 2022; Shao et al., 2024; Kim et al., 2024). Specifically, we use 32 sequences for OPT models and 128 sequences for LLaMA models. Each sequence consists of 2,048 tokens, sampled from the first shard of the C4 dataset.
Latency Profiling. Using Torch CUDA profiler, we measure single-sequence (batch size 1) generation of 1024 tokens on a single NVIDIA RTX 4090 GPU, reporting CUDA time and peak memory usage with LUT-based inference kernels in (Kim et al., 2024).